Shift report 20000808 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 08 2000 - 15:47:25 EDT

  • Next message: http account: "Shift report 20000809 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Aug  8 15:47:23 2000, Shift: 08:00-16:00
    Supervisor: D Beavis, with: Erik Jakobsen and Claus Jorgensen
    Log book#4 pages 30 - 31
    
    Shift Summary:
    ------------------------------------------------------------
    NO stored beam all shift.
    Many inject attempts and some acc. attempts.
    TPC anodes have been off all shift.
    Spent the shift trying to resolve various TOFW bad channels.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 08 2000 - 15:53:02 EDT